# {tocentry}[Uncaptioned image]

### v3

@@ Abstract
Quantum Selected Configuration Interaction (QSCI) methods (also known as Sample-based Quantum Diagonalization, SQD) have emerged as promising near-term approaches to solving the electronic Schrödinger equation with quantum computers. In this work, we perform numerical analysis to show that QSCI methods face critical limitations that severely hinder their practical applicability in chemistry. Using the nitrogen molecule and the iron-sulfur cluster [2Fe-2S] as examples, we demonstrate that while QSCI can, in principle, yield high-quality CI expansions similar to classical SCI heuristics in some cases, the method struggles with inefficiencies in finding new determinants as sampling repeatedly selects already seen configurations. This inefficiency becomes especially pronounced when targeting high-accuracy results or sampling from an approximate ansatz. In cases where the sampling problem is not present, the resulting CI expansions are less compact than those generated from classical heuristics, rendering QSCI an overall more expensive method. Our findings suggest a significant drawback in QSCI methods when sampling from the ground-state distribution as the inescapable trade-off between finding sufficiently many determinants and generating compact, accurate CI expansions. This ultimately hinders utility in quantum chemistry applications, as QSCI falls behind more efficient classical counterparts.
@@ 1 Introduction
The full configuration interaction (FCI) method provides the exact solution to the electronic Schrödinger equation (within a given basis) and is thus considered the ground truth within quantum chemistry1, 2, 3. However, due to the exponential scaling in computing and memory requirements, exact FCI is limited to very small molecules. One of the largest demonstrations to date was by Gao et al. on propane (C3H8)/STO-3G4, who used a distributed FCI implementation on 256 nodes to treat a CI expansion of 1.3 trillion determinants. As a technical achievement, such work is certainly impressive. If one is willing to settle for near-exact FCI results (within, e.g., micro-Hartree precision), approximate methods such as selected configuration interaction (SCI) can obtain results much more affordably5. Various developments within electronic structure theory have focused on achieving near-FCI accuracy at a lower cost3, including methods based on the density matrix renormalization group (DMRG)6, many-body expanded FCI (MBE-FCI)7, 8, 9, and various flavors of SCI10, 11, 12, 13. Recently, such methods were put to the test in a “blind challenge” of predicting the ground-state energy of the benzene molecule (C6H6) in a cc-pVDZ basis set2. It revealed that a firm consensus between various near-exact FCI methods has still not been reached, even when agreement from calculations on smaller systems would have suggested so.
@@
Applying quantum computers within the field of quantum chemistry has received significant attention in recent years14, 15, 16, 17. This is partly due to the expectation that quantum computers should be a natural candidate for solving the electronic Schrödinger equation as they inherently operate within an exponentially large Hilbert space. In the long term, provided fault-tolerant devices become a reality, quantum computers could, for example, be used to obtain ground state energies using quantum phase estimation18, 19. However, current near-term quantum computers are noisy and rather limited in terms of qubit count, error rates, and execution speed, which limits their practical advantage over classical computing architectures20. This has motivated the development of hybrid algorithms that take advantage of both quantum and classical computing architectures, resulting in shallower quantum circuits, which are more realistic options for practical quantum computing applications in the near term21, 22, 23. The variational quantum eigensolver (VQE) algorithm24, 25, 26 has emerged as a particularly popular near-term method. VQE-type methods rely on a parameterized ansatz, which is iteratively optimized using a hybrid quantum-classical feedback loop. The quantum device prepares a state and measures the expectation values over Pauli operators. The classical computer collects these measurements to evaluate expectation values over the fermionic Hamiltonian and updates the parameters of the parameterized ansatz using classical optimization methods. Although popular, VQE-type methods have several serious drawbacks related to the measurement process,27 which inherently has statistical noise, and the expressibility of the ansatz28, 29, 26.
@@
The quantum selected configuration interaction (QSCI) method, introduced by Kanno et al.30, 31, addresses some of the challenges faced with VQE by drastically reducing the quantum workload, leveraging quantum devices solely for determinant selection. The central idea of the method is that some approximate ground state is prepared on a quantum computer, which is then used to draw samples corresponding to determinants in the computational basis. The CI Hamiltonian is then classically diagonalized in the basis of the selected determinants using conventional SCI diagonalization routines. We note that the classic diagonalization eventually becomes the computational bottleneck of the algorithm, provided enough determinants are discovered. Thus, the “quantum” part of QSCI lies solely in the selection of determinants for the SCI expansion. Such a division of work between quantum and classical resources has some clear advantages. First, since the quantum device is only used for selection, while energies are obtained classically, one removes the statistical errors associated with Hamiltonian measurement. Second, the CI vector is accessible on the classical device after the diagonalization, which allows computing expectation values over other quantum mechanical operators or correcting the energy of the variational wave function with Epstein-Nesbet perturbation theory32. In the original QSCI work, the dimension of the subspace (i.e., the number of determinants included in the CI expansion) was rather limited, with at most around a hundred determinants, in part due to a limited variational space when using rather small basis sets.
@@
The QSCI method has also been demonstrated on a larger scale by Robledo-Moreno et al.33, under the name sample-based quantum diagonalization (SQD) with applications to potential energy curves of the nitrogen molecule (in a cc-pVDZ basis) and the iron-sulfur cluster model systems [2Fe-2S] and [4Fe-4S], with active spaces of (30e,20o) and (54e,36o), respectively. The demonstrations in Ref. 33 are relatively “large-scale” in nature, using CI expansion with millions of determinants. The central procedure is similar to the original QSCI paper but demonstrates the use of a large number of qubits (up to 77) and handling aspects such as treating quantum samples with incorrect particle numbers. The state preparation is handled with a local unitary cluster Jastrow ansatz (LUCJ)34, 35, with parameters assigned classically based on CCSD $t_{2}$-amplitudes. The SQD approach has since become a showcase for quantum-centric supercomputing36 and has been applied in several follow-up works, including the treatment of interaction energies with the potential energy surfaces of water and methane dimers37, spin state energetics of Fe(III)-NTA 38 and methylene39, the treatment of excited states40, inclusion of environment effects through density matrix embedding theory41, in combination with implicit-solvation models42, and combined with phaseless auxiliary- field quantum Monte Carlo43.
@@
Separately, we note that an approach for determining the ground state of the SCI using quantum computers has also been proposed44. Unlike QSCI and SQD, determinants are selected using conventional SCI heuristics, and the focus is on leveraging variational quantum circuits to find the ground-state energy of the CI Hamiltonian.
@@
The QSCI method introduces a potentially useful approach to determinant selection via quantum sampling, but several fundamental challenges limit its practical utility in real-world applications. First of all, the method arguably requires the preparation of a good state on the quantum device. Ideally, such a state should provide a probability distribution that closely matches the exact FCI solution since one would otherwise miss important determinants or include a significant fraction of irrelevant ones. Preparing such states is difficult, and it remains unclear how such a state preparation could take place without solving the electronic Schrödinger equation to high accuracy in the first place. Independently, two groups recently proposed methods with state-preparation protocols based on time-evolved or Hamiltonian-simulation QSCI, which could perhaps provide a workaround for this problem45, 46. Later, similar developments under the SQD moniker have independently appeared47.
@@
A second issue arises from the noisy nature of currently available quantum devices. Even though the target ansatz is, in principle, particle-conserving (i.e., has a desired number of $\alpha$ and $\beta$ electrons), the presence of hardware noise leads to many samples having the wrong number of particles. For example, in Ref. 33, only 0.44% of samples for the iron-sulfur cluster [2Fe-2S] (30e,20o) are within the target particle sector. Solutions to this problem have been proposed in the form of methods like the self-consistent configuration recovery (S-CORE)33, which seek to recover nearby bit-strings with a correct particle number from the set of erroneous bitstrings.
@@
Ultimately, we demonstrate that even with an ideal sampling wave function in a noise-free environment, the stochastic nature of sampling itself either introduces significant inefficiencies or leads to less compact wave functions. Less compact CI expansions (i.e., more determinants for the same energy) lead to higher (classical) computational costs associated with the diagonalization of the CI Hamiltonian for the same final accuracy in the computation. These limitations undermine any practical advantage of QSCI over well-established classical SCI heuristics. Furthermore, enhancements like ext-SQD40, which add determinants post-sampling, could equally be applied to classical SCI methods, negating any unique benefit of quantum sampling.
@@ 2 Theory
with the CI coefficients, $c_{I}$. The included determinants can be selected by excitation level out of a reference determinant (e.g., Hartree-Fock), forming the CISD, CISDT, etc. hierarchy48, which eventually terminates at the FCI expansion when $N$-tuple excitations are included. Alternatively, as is done in SCI expansions10, 11, 12, 13, a smaller set of important determinants can be selected based on classical heuristics. This approach recognizes that only a small fraction of determinants in the FCI expansion contribute significantly to the electronic wave function for most molecular systems 49.
@@
In addition to QSCI, we will also consider a representative of classical SCI. In particular, we will consider Heat-bath Configuration Interaction12 (HCI). In HCI, important determinants for the CI expansion are added based on their expected contribution to the wave function. It uses a parameter $\varepsilon$ to screen a potential determinant $\left|\Phi_{I}\right\rangle$ and includes it if any of the existing determinants in the SCI expansion satisfy
@@
The HCI method works in an iterative fashion. Starting from some initial CI expansion (e.g., the HF determinant), the following steps are repeated: 1) the CI Hamiltonian is diagonalized, and the CI coefficients of the current determinant set are obtained; 2) determinants are added according to the criterion in eq. (3), with $\varepsilon$ as a parameter. These steps are repeated until no more determinants can be added according to eq. (3) (in the original work12, when the number of new determinants is less than 1% of the number of determinants already selected).
@@ 3 Computational Details
Hartree-Fock and CASCI (10e, 22o) wave functions, as well as one- and two-electron molecular orbital (MO) integrals (cc-pVDZ50 basis) of the nitrogen molecule, were obtained using PySCF 51. We considered structures near the equilibrium geometry ($R=1.09$Å) and along the potential energy surface $R=0.7$Å to $R=3.0$Å. Additionally, we investigated the iron-sulfur cluster [2Fe-2S] (30e,20o), where the one- and two-electron integrals were taken from the FCIDUMP file published by Li and Chan52, 53. Finally, we also include results to assess the scalability of QSCI with system size, using linear lithium hydride chains, $(\mathrm{LiH})_{n}$/STO-3G, as an example, applying full-valence CAS($2n$e, $5n$o) active spaces. For this, we consider a weakly correlated regime $R_{\mathrm{Li-H}}=1.61$, $R_{\mathrm{LiH-LiH}}=1.81$, and a strongly correlated regime with $R_{\mathrm{Li-H}}=3.22$, $R_{\mathrm{LiH-LiH}}=3.62$. Throughout the paper, we adopted the notation ($N$e, $M$o) for the active spaces, where $N$ is the number of active electrons and $M$ is the number of active orbitals.
The QSCI calculations were performed by drawing determinant samples from the CI vectors from the CASCI calculations with the probability of selecting a particular determinant given as $|c_{I}|^{2}$. The number of samples corresponds to the number of times the ansatz is measured. Our approach constitutes an ideal limit of QSCI since it assumes perfect state preparation of the exact CI vector and avoids any quantum device noise (i.e., non-ideal execution of the quantum circuits). The PyCI library54 was used to diagonalize the resulting CI Hamiltonian. HCI calculations were also performed using PyCI.
To assess the impact of an approximate ansatz in the sampling routine of QSCI, we obtained state vectors of equilibrium N2 for the unitary cluster Jastrow (UCJ)34 and local UCJ (LUCJ)35 ansatz using ffsim55. For both methods, two layers were used, and the ansatz parameters were assigned using the $t_{2}$ amplitudes of a classic CCSD simulation. This mimics the state preparation procedure used in recent works of IBM Quantum employing SQD/QSCI on quantum hardware33.
@@ 4.1 Potential energy curves of the nitrogen molecule
Another way to see the severity of the repeated sampling problem is to compare the search costs of QSCI and HCI when increasing the size (i.e., accuracy) of the CI expansion. As shown in Figure S4, the problem of repeatedly finding the same determinants with QSCI manifests as a super-linear (nearly quadratic) search cost scaling for N2. In contrast, HCI shows only a linear increase in the search cost as a function of the number of determinants in the CI expansion. The HCI method progressively grows the CI expansion through successive iterations. In addition to determinant selection, this requires repeated diagonalizations of the SCI Hamiltonian. Thus, with HCI, one should also factor in the cumulative costs of diagonalization. In contrast, QSCI in an ideal setting requires only a single round of diagonalization. Even when considering the total calculation costs of HCI, including determinant search, sparse Hamiltonian construction, and Davidson diagonalization, the observed scaling remains more modest for HCI (see Figures S4 and S5).

### v4

@@ Abstract
Quantum Selected Configuration Interaction (QSCI) methods (also known as Sample-based Quantum Diagonalization, SQD) have emerged as promising near-term approaches to solving the electronic Schrödinger equation with quantum computers. In this work, we perform numerical analysis to show that QSCI methods face critical limitations that severely hinder their practical applicability in chemistry. Using the nitrogen molecule and the iron-sulfur cluster [2Fe-2S] as examples, we demonstrate that while QSCI can, in principle, yield high-quality configuration interaction (CI) expansions similar to classical SCI heuristics in some cases, the method struggles with inefficiencies in finding new determinants as sampling repeatedly selects already seen configurations. This inefficiency becomes especially pronounced when targeting high-accuracy results or sampling from an approximate ansatz. In cases where the sampling problem is not present, the resulting CI expansions are less compact than those generated from classical heuristics, rendering QSCI an overall more expensive method. Our findings suggest a significant drawback in QSCI methods when sampling from the ground-state distribution as the inescapable trade-off between finding sufficiently many determinants and generating compact, accurate CI expansions. This ultimately hinders utility in quantum chemistry applications, as QSCI falls behind more efficient classical counterparts.
@@ 1 Introduction
The full configuration interaction (FCI) method provides the exact solution to the electronic Schrödinger equation (within a given basis) and is thus considered the ground truth within quantum chemistry1, 2, 3. However, due to the exponential scaling in computing and memory requirements, exact FCI is limited to very small molecules. One of the largest demonstrations to date was by Gao et al. on propane (C3H8)/STO-3G4, who used a distributed FCI implementation on 256 nodes to treat a CI expansion of 1.3 trillion determinants. Recent work using a small-tensor-product distributed active space representation has even reported calculations exceeding a quadrillion determinants 5. As a technical achievement, such work is certainly impressive. If one is willing to settle for near-exact FCI results (within, e.g., micro-Hartree precision), approximate methods such as selected configuration interaction (SCI) can obtain results much more affordably6. Various developments within electronic structure theory have focused on achieving near-FCI accuracy at a lower cost3, including methods based on the density matrix renormalization group (DMRG)7, many-body expanded FCI (MBE-FCI)8, 9, 10, and various flavors of SCI11, 12, 13, 14. Recently, such methods were put to the test in a “blind challenge” of predicting the ground-state energy of the benzene molecule (C6H6) in a cc-pVDZ basis set2. It revealed that a firm consensus between various near-exact FCI methods has still not been reached, even when agreement from calculations on smaller systems would have suggested so.
@@
Applying quantum computers within the field of quantum chemistry has received significant attention in recent years15, 16, 17, 18. This is partly due to the expectation that quantum computers should be a natural candidate for solving the electronic Schrödinger equation as they inherently operate within an exponentially large Hilbert space. In the long term, provided fault-tolerant devices become a reality, quantum computers could, for example, be used to obtain ground state energies using quantum phase estimation19, 20. However, current near-term quantum computers are noisy and rather limited in terms of qubit count, error rates, and execution speed, which limits their practical advantage over classical computing architectures21. This has motivated the development of hybrid algorithms that take advantage of both quantum and classical computing architectures, resulting in shallower quantum circuits, which are more realistic options for practical quantum computing applications in the near term22, 23, 24. The variational quantum eigensolver (VQE) algorithm25, 26, 27 has emerged as a particularly popular near-term method. VQE-type methods rely on a parameterized ansatz, which is iteratively optimized using a hybrid quantum-classical feedback loop. The quantum device prepares a state and measures the expectation values over Pauli operators. The classical computer collects these measurements to evaluate expectation values over the fermionic Hamiltonian and updates the parameters of the parameterized ansatz using classical optimization methods. Although popular, VQE-type methods have several serious drawbacks related to the measurement process,28 which inherently has statistical noise, and the expressibility of the ansatz29, 30, 27.
@@
The quantum selected configuration interaction (QSCI) method, introduced by Kanno et al.31, 32, addresses some of the challenges faced with VQE by drastically reducing the quantum workload, leveraging quantum devices solely for determinant selection. The central idea of the method is that some approximate ground state is prepared on a quantum computer, which is then used to draw samples corresponding to determinants in the computational basis. The CI Hamiltonian is then classically diagonalized in the basis of the selected determinants using conventional SCI diagonalization routines. We note that the classic diagonalization eventually becomes the computational bottleneck of the algorithm, provided enough determinants are discovered. Thus, the “quantum” part of QSCI lies solely in the selection of determinants for the SCI expansion. Such a division of work between quantum and classical resources has some clear advantages. First, since the quantum device is only used for selection, while energies are obtained classically, one removes the statistical errors associated with Hamiltonian measurement. Second, the CI vector is accessible on the classical device after the diagonalization, which allows computing expectation values over other quantum mechanical operators or correcting the energy of the variational wave function with Epstein-Nesbet perturbation theory33. In the original QSCI work, the dimension of the subspace (i.e., the number of determinants included in the CI expansion) was rather limited, with at most around a hundred determinants, in part due to a limited variational space when using rather small basis sets.
@@
The QSCI method has also been demonstrated on a larger scale by Robledo-Moreno et al.34, under the name sample-based quantum diagonalization (SQD) with applications to potential energy curves of the nitrogen molecule (in a cc-pVDZ basis) and the iron-sulfur cluster model systems [2Fe-2S] and [4Fe-4S], with active spaces of (30e,20o) and (54e,36o), respectively. The demonstrations in Ref. 34 are relatively “large-scale” in nature, using CI expansion with millions of determinants. The central procedure is similar to the original QSCI paper but demonstrates the use of a large number of qubits (up to 77) and handling aspects such as treating quantum samples with incorrect particle numbers. The state preparation is handled with a local unitary cluster Jastrow ansatz (LUCJ)35, 36, with parameters assigned classically based on CCSD $t_{2}$-amplitudes. The SQD approach has since become a showcase for quantum-centric supercomputing37 and has been applied in several follow-up works, including the treatment of interaction energies with the potential energy surfaces of water and methane dimers38, spin state energetics of Fe(III)-NTA 39 and methylene40, the treatment of excited states41, inclusion of environment effects through density matrix embedding theory42, in combination with implicit-solvation models43, and combined with phaseless auxiliary- field quantum Monte Carlo44.
@@
Separately, we note that an approach for determining the ground state of the SCI using quantum computers has also been proposed45. Unlike QSCI and SQD, determinants are selected using conventional SCI heuristics, and the focus is on leveraging variational quantum circuits to find the ground-state energy of the CI Hamiltonian.
@@
The QSCI method introduces a potentially useful approach to determinant selection via quantum sampling, but several fundamental challenges limit its practical utility in real-world applications. First of all, the method arguably requires the preparation of a good state on the quantum device. Ideally, such a state should provide a probability distribution that closely matches the exact FCI solution since one would otherwise miss important determinants or include a significant fraction of irrelevant ones. Preparing such states is difficult, and it remains unclear how such a state preparation could take place without solving the electronic Schrödinger equation to high accuracy in the first place. Independently, two groups recently proposed methods with state-preparation protocols based on time-evolved or Hamiltonian-simulation QSCI, which could perhaps provide a workaround for this problem46, 47. Later, similar developments under the SQD moniker have independently appeared48.
@@
A second issue arises from the noisy nature of currently available quantum devices. Even though the target ansatz is, in principle, particle-conserving (i.e., has a desired number of $\alpha$ and $\beta$ electrons), the presence of hardware noise leads to many samples having the wrong number of particles. For example, in Ref. 34, only 0.44% of samples for the iron-sulfur cluster [2Fe-2S] (30e,20o) are within the target particle sector. Solutions to this problem have been proposed in the form of methods like the self-consistent configuration recovery (S-CORE)34, which seek to recover nearby bit-strings with a correct particle number from the set of erroneous bitstrings.
@@
Ultimately, we demonstrate that even with an ideal sampling wave function in a noise-free environment, the stochastic nature of sampling itself either introduces significant inefficiencies or leads to less compact wave functions. Less compact CI expansions (i.e., more determinants for the same energy) lead to higher (classical) computational costs associated with the diagonalization of the CI Hamiltonian for the same final accuracy in the computation. These limitations undermine any practical advantage of QSCI over well-established classical SCI heuristics. Furthermore, enhancements like ext-SQD41, which add determinants post-sampling, could equally be applied to classical SCI methods, negating any unique benefit of quantum sampling.
@@ 2 Theory
with the CI coefficients, $c_{I}$. The included determinants can be selected by excitation level out of a reference determinant (e.g., Hartree-Fock), forming the CISD, CISDT, etc. hierarchy49, which eventually terminates at the FCI expansion when $N$-tuple excitations are included. Alternatively, as is done in SCI expansions11, 12, 13, 14, a smaller set of important determinants can be selected based on classical heuristics. This approach recognizes that only a small fraction of determinants in the FCI expansion contribute significantly to the electronic wave function for most molecular systems 50.
@@
In addition to QSCI, we will also consider a representative of classical SCI. In particular, we will consider Heat-bath Configuration Interaction13 (HCI). In HCI, important determinants for the CI expansion are added based on their expected contribution to the wave function. It uses a parameter $\varepsilon$ to screen a potential determinant $\left|\Phi_{I}\right\rangle$ and includes it if any of the existing determinants in the SCI expansion satisfy
@@
The HCI method works in an iterative fashion. Starting from some initial CI expansion (e.g., the HF determinant), the following steps are repeated: 1) the CI Hamiltonian is diagonalized, and the CI coefficients of the current determinant set are obtained; 2) determinants are added according to the criterion in eq. (3), with $\varepsilon$ as a parameter. These steps are repeated until no more determinants can be added according to eq. (3) (in the original work13, when the number of new determinants is less than 1% of the number of determinants already selected).
@@ 3 Computational Details
Hartree-Fock and CASCI (10e, 22o) wave functions, as well as one- and two-electron molecular orbital (MO) integrals (cc-pVDZ51 basis) of the nitrogen molecule, were obtained using PySCF 52. We considered structures near the equilibrium geometry ($R=1.09$Å) and along the potential energy surface $R=0.7$Å to $R=3.0$Å. Additionally, we investigated the iron-sulfur cluster [2Fe-2S] (30e,20o), where the one- and two-electron integrals were taken from the FCIDUMP file published by Li and Chan53, 54. Finally, we also include results to assess the scalability of QSCI with system size, using linear lithium hydride chains, $(\mathrm{LiH})_{n}$/STO-3G, as an example, applying full-valence CAS($2n$e, $5n$o) active spaces. For this, we consider a weakly correlated regime $R_{\mathrm{Li-H}}=1.61$, $R_{\mathrm{LiH-LiH}}=1.81$, and a strongly correlated regime with $R_{\mathrm{Li-H}}=3.22$, $R_{\mathrm{LiH-LiH}}=3.62$. Throughout the paper, we adopted the notation ($N$e, $M$o) for the active spaces, where $N$ is the number of active electrons and $M$ is the number of active orbitals.
The QSCI calculations were performed by drawing determinant samples from the CI vectors from the CASCI calculations with the probability of selecting a particular determinant given as $|c_{I}|^{2}$. The number of samples corresponds to the number of times the ansatz is measured. Our approach constitutes an ideal limit of QSCI since it assumes perfect state preparation of the exact CI vector and avoids any quantum device noise (i.e., non-ideal execution of the quantum circuits). The PyCI library55 was used to diagonalize the resulting CI Hamiltonian. HCI calculations were also performed using PyCI.
To assess the impact of an approximate ansatz in the sampling routine of QSCI, we obtained state vectors of equilibrium N2 for the unitary cluster Jastrow (UCJ)35 and local UCJ (LUCJ)36 ansatz using ffsim56. For both methods, two layers were used, and the ansatz parameters were assigned using the $t_{2}$ amplitudes of a classic CCSD simulation. This mimics the state preparation procedure used in recent works of IBM Quantum employing SQD/QSCI on quantum hardware34.
@@ 4.1 Potential energy curves of the nitrogen molecule
Another way to see the severity of the repeated sampling problem is to compare the search costs of QSCI and HCI when increasing the size (i.e., accuracy) of the CI expansion. As shown in Figure S4, the problem of repeatedly finding the same determinants with QSCI manifests as a super-linear (nearly quadratic) search cost scaling for N2. In contrast, HCI shows only a linear increase in the search cost as a function of the number of determinants in the CI expansion. The HCI method progressively grows the CI expansion through successive iterations. In addition to determinant selection, this requires repeated diagonalizations of the SCI Hamiltonian. Thus, with HCI, one should also factor in the cumulative costs of diagonalization. In contrast, QSCI in an ideal setting requires only a single round of diagonalization. Even when considering the total calculation costs of HCI, including determinant search, sparse Hamiltonian construction, and Davidson diagonalization, the observed scaling in determinants of HCI remains more modest at around $N_{\mathrm{det}}^{1.2}-N_{\mathrm{det}}^{1.3}$ for the total calculation time than just the search time of QSCI (excluding costs associated with diagonalization) at around $N_{\mathrm{det}}^{2.2}-N_{\mathrm{det}}^{2.9}$ (see Figures S4 and S5 and the surrounding discussion). A direct comparison of computational timings requires assigning a conversion factor from quantum samples to seconds, which we will consider in Section 4.4.